OpenStudy (kimberly_pr):

Graph the first six terms of a sequence where a1 = 4 and r = 2

OpenStudy (kimberly_pr):

@Atsie @Qwertty123 @dumbcow

OpenStudy (welshfella):

this is a geometric sequence a1 , a1*r, a1*r^2, a1*r^3 .... a1*r^5

OpenStudy (welshfella):

so first 2 terms are 4 , 4*2

OpenStudy (kimberly_pr):

4*2= 8 So what next? Do I just plug in 4 to all the equations you gave me? @welshfella

OpenStudy (welshfella):

yes the next term is 4 * 2^2

OpenStudy (welshfella):

then 4 * 2^3 do you see the pattern?

OpenStudy (kimberly_pr):

Next would be 4*2^2= 16

OpenStudy (welshfella):

right

OpenStudy (kimberly_pr):

So I think the answer would be the graph with points (1,8) (2,16) (3,32) (4,64) (5,128) (6, 256) right?

OpenStudy (kimberly_pr):

Or am I wrong?... I'm not so sure anymore but I THINK ^^ those points are right

OpenStudy (welshfella):

yes - sorry i mad a mistake - you multiply by 2 not by 4 Your results are correct

OpenStudy (kimberly_pr):

Thank you!! :)

OpenStudy (welshfella):

yw

OpenStudy (kimberly_pr):

Can you help me with one more?

OpenStudy (welshfella):

I have 5 minutes only bu t ok

OpenStudy (kimberly_pr):

Thanks Calculate the average rate of change for the graphed sequence from n = 2 to n = 4.

OpenStudy (kimberly_pr):

And the points on the graph are (1,-3) (2,-3.5) (3,-6.75) (4,-10.25) (5,-15.1875) (6,-22.78125)

OpenStudy (welshfella):

average rate of change = 4th term - 2d term ----------------- 4 - 2 = 64 - 16 ------ = 48 / 2 4 - 2

OpenStudy (welshfella):

gotta go

OpenStudy (kimberly_pr):

@Qwertty123 @Atsie

OpenStudy (kimberly_pr):

@dinamix

OpenStudy (kimberly_pr):

@phi

OpenStudy (kimberly_pr):

@dinamix and @phi I need help on my second question, not my first one

OpenStudy (phi):

can you make it a new post. It's tough to follow this.

OpenStudy (kimberly_pr):

Ok, I will :)